# Geometry and cohomology of compactified Deligne–Lusztig varieties

### v2

@@ 4.1. $\mathbb{P}^{1}$-fibrations over Deligne–Lusztig varieties
Let $w_{1},\ldots,w_{r},w_{1}^{\prime},\ldots,w_{m}^{\prime}\in W$ and $s\in S$. Then the projection map
@@
given by
$(B_{0},\ldots,B_{r},B_{r+1},B_{r+2},\ldots,B_{r+m+1},FB_{0})\mapsto(B_{0},\ldots,B_{r},B_{r+2},\ldots,B_{r+m+1},FB_{0})$ is a $G^{F}$-equivariant Zariski $\mathbb{P}^{1}$-bundle.
If $F(s)\in S$, then the $G^{F}$-equivariant morphism of $k$-schemes
@@
For any $t\in S$, the projection map $O(\underline{t},\underline{t})\to O(\underline{t})$ given by $(B_{0},B_{1},B_{2})\mapsto(B_{0},B_{2})$ is a $\mathbb{P}^{1}$-bundle. Recall that it is a property of the Bruhat decomposition that if $(B_{0},B_{1})\in O(t)$ and $(B_{1},B_{2})\in O(t)$ then $(B_{0},B_{2})\in O(\underline{t})$, cf. [DL, §1.2].
@@
Assume $F(s)\in S$ and consider $t=F(s)$. One can verify that $\pi$ is $G^{F}$-equivariant by its definition. Let $(B_{0},B_{1})\in O(\underline{s})$, so $(FB_{0},FB_{1})\in O(\underline{F(s)})$. Recall that via the projection to the first $(r+2)$-entries, $X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$ can be identified with a closed subscheme of $(G/B)^{r+2}$. Under this identification, we have an embedding
@@
Let $w_{1},\ldots,w_{r}\in W$ and $s\in S$ such that $F(s)\in S$ and $\overline{Bw_{n}B}$ is smooth for $n=1,\ldots,r$. The morphisms
@@
Given $(B_{0},B_{1})\in O(\underline{s})$, we know that $(FB_{0},FB_{1})\in O(F(\underline{s}))$. By construction we have $(B_{r+1},FB_{0})\in O(F(\underline{s}))$, so $(B_{r+1},FB_{1})\in O(F(\underline{s}))$ and $(FB_{r+1},F^{2}B_{1})\in O(F(\underline{s}))$ [DL, §1.2]. Thus $\tau_{1}$ and $\tau_{2}$ are well-defined. They are $G^{F}$-equivariant by construction.
@@
${Z}$${O(\underline{F^{-1}(s)})}$${X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})}$${G/B,}$$\scriptstyle{r_{2}}$$\scriptstyle{d_{1}}$$\scriptstyle{{\rm pr}_{1}}$$\scriptstyle{{\rm pr}_{0}}$ (4)
@@
and that $\iota$ is surjective. By the Bruhat decomposition, for every $i=1,\ldots,r$, there exists $v_{i}\in W$ such that $(B_{i},B_{i}^{\prime})\in O(v_{i})$. This implies $(FB_{i},FB_{i}^{\prime})\in O(F(v_{i}))$. We know $FB_{i}=FB_{i}^{\prime}$ by assumption, so $F(v_{i})=e$ is the neutral element in $W$ and $v_{i}=e$ because $F:W\to W$ is an automorphism, cf. [DL, §1.2]. It follows that $\iota$ is also injective and an isomorphism.
Hence $d_{1}$ is a Zariski $\mathbb{P}^{1}$-bundle. Since $F:X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})\to X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})$ is faithfully flat by our assumption on smoothness [GoerWedhAlgGeom1, Corollary 14.129], we deduce that $\tau_{2}$ is faithfully flat by faithfully flat descent [GoerWedhAlgGeom1, Corollary 14.12, Proposition 14.3(2)].
@@
Let $w_{1},\ldots,w_{r}\in W$ and $s\in S$ such that $F(s)\in S$ and $\overline{Bw_{n}B}$ is smooth for $n=1,\ldots,r$. There are $G^{F}$-equivariant isomorphisms for all $i\geq 0$,
@@
The proof is simpler for the Zariski $\mathbb{P}^{1}$-bundles $\phi$ and $\pi$, so we only record the proof for the fppf $\mathbb{P}^{1}$-bundles $\tau_{1}$ and $\tau_{2}$. We use the notations of Lemma 4.1.3. Since $d_{1}$ is a Zariski $\mathbb{P}^{1}$-bundle, it is proper and faithfully flat. Hence $\tau_{1}$ and $\tau_{2}$ are proper and faithfully flat by faithfully flat descent [GoerWedhAlgGeom1, Proposition 14.53].
@@
Let $x\in X_{2}$ and $y\in X_{2}^{\prime}$ such that $F(y)=x$ as in diagram (3). By Lemma 4.1.3, the base change of the fibre $x\times_{X_{2}}X_{1}$ to ${\rm Spec}(\kappa(y))$ is isomorphic to $\mathbb{P}^{1}_{\kappa(y)}$. It follows from flat base change that $H^{j}(x\times_{X_{2}}X_{1},\mathcal{O}_{x\times_{X_{2}}X_{1}})=0$ for all $j\geq 0$ [GoerWedhAlgGeom2, Corollary 22.91]. We conclude by Grauert’s theorem [GoerWedhAlgGeom2, Proposition 23.142] (cf. [GoerWedhAlgGeom1, Proposition 3.35]) that $R^{j}f_{*}\mathcal{O}_{X_{1}}=0$ for all $j>0$. Hence the Leray spectral sequence degenerates.
We have an fppf covering $\mathfrak{U}=\{\gamma_{\alpha}:U_{\alpha}\to X_{2}\}_{\alpha}$ of $X_{2}$ in Lemma 4.1.3, such that the base change $f_{\alpha}:X_{1}\times_{X_{2}}U_{\alpha}\to U_{\alpha}$ of $f$ along $\gamma_{\alpha}$ is a trivial $\mathbb{P}^{1}$-bundle for all $\alpha$. Since $U_{\alpha}$ is affine and $\gamma_{\alpha}$ is an affine morphism for all $\alpha$ [GoerWedhAlgGeom1, Proposition 12.3(3)], any simplex $V$ in the Čech nerve of $\mathfrak{U}$ is affine and for any coherent sheaf $\mathcal{F}$ on $X_{2}$, we have $H^{i}(X_{2},\mathcal{F})\overset{\sim}{\longrightarrow}\check{H}^{i}(\mathfrak{U},\mathcal{F})$ for all $i\geq 0$ [Godement, Theorem II.5.4.1, Corollary]. Note that $f_{*}\mathcal{O}_{X_{1}}$ is a coherent $\mathcal{O}_{X_{2}}$-module because $f$ is proper and cohomology of coherent sheaves is independent of topology [stacks-project, 03DW].
For each simplex $V$ in the Čech nerve of $\mathfrak{U}$, we have $V\times_{X_{2}}X_{1}\overset{\sim}{\longrightarrow}V\times_{{\rm Spec}\ k}\mathbb{P}^{1}$, so $f_{*}\mathcal{O}_{X_{1}}(V)=\mathcal{O}_{V}(V)$ by flat base change [GoerWedhAlgGeom2, Proposition 22.90]. This identification is compatible with restriction of sections, so the Čech complexes on $\mathfrak{U}$ for $f_{*}\mathcal{O}_{X_{1}}$ and $\mathcal{O}_{X_{2}}$ are identical and $H^{i}(X_{2},\mathcal{O}_{X_{2}})\overset{\sim}{\longrightarrow}H^{i}(X_{2},f_{*}\mathcal{O}_{X_{1}})$ for all $i\geq 0$.
@@
$H^{i}\left(X_{2},\mathcal{O}_{X_{1}}\right)\overset{\sim}{\longrightarrow}H^{i}\left(X_{1},\mathcal{O}_{X_{1}}\right),$
@@
The $G^{F}$-equivariant isomorphisms for the cohomology of $\omega_{X_{1}}$ and $\omega_{X_{2}}$ follow from Serre duality for Deligne–Mumford stacks [Nironi, Corollary 2.10].
@@
Let $s\in S$ and assume $F(s)\in S$. Let $w=sv$ and $w^{\prime}=vF(s)$ be elements of $S^{*}$. There are $G^{F}$-equivariant isomorphisms for all $i\geq 0$,
@@ 4.2. $F$-conjugacy classes in $W$
Minimal length elements were studied by Geck, Kim, and Pfeiffer [GeckKimPfeifferMinlength], as well as He and Nie [HeNieminlength]. Under the assumption $F(S)=S$, these two papers showed that for any $w\in W$, there exists a minimal length element $w^{\prime}$ in the $F$-conjugacy class of $w$ such that $w\to_{F}w^{\prime}$ (cf. Definition 4.2.1). In combination with Proposition 4.1.4, this result gives an inductive approach for studying cohomology groups of Deligne–Lusztig varieties.
@@
Furthermore, for any pairs of minimal length elements $w$ and $w^{\prime}$ in an elliptic $F$-conjugacy class $C$ in $W$, we have $w\to_{F}w^{\prime}$ and $w^{\prime}\to_{F}w$, cf. [HeNieminlength, Corollary 4.4]. The definition of an elliptic $F$-conjugacy class, which we do not recall here, can be found in §4.1 of loc. cit.
@@ 5.1. Main theorem and applications
Recall that the braid monoid associated to $W$ is a monoid generated by $S$ and all the relations of the form $P(s,t)=P(t,s)$ with $s,t\in S$, $t\neq s$, such that $(st)^{m}=1$ is a braid relation in $W$, cf. [GeckPfeiffer, §4]. We denote the braid monoid by $B^{+}$. There is a natural surjective morphism of monoids $\beta:S^{*}\twoheadrightarrow B^{+}$.
@@
Assume $F(S)=S$. For each $w\in S^{*}$, there exists an $F$-conjugacy class $C$ in $W$ and a minimal length element $x_{0}\in C$ such that for any reduced expression $x\in S^{*}$ of $x_{0}$, there are $G^{F}$-equivariant isomorphisms
@@
We prove the theorem for the structure sheaf. The results for the canonical sheaf follows from Serre duality for Deligne–Mumford stacks [Nironi, Corollary 2.10].
Let $w\in S^{*}$. Consider the image $\beta(w)\in B^{+}$ of $w$ under the projection $\beta:S^{*}\twoheadrightarrow B^{+}$ and let $u\in S^{*}$ be another lift of $\beta(w)$, then we have
$H^{i}\Big{(}{\overline{X}^{S^{*}}(u)},\mathcal{O}_{{\overline{X}^{S^{*}}(u)}}\Big{)}\overset{\sim}{\longrightarrow}H^{i}\Big{(}{\overline{X}^{S^{*}}(w)},\mathcal{O}_{{\overline{X}^{S^{*}}(w)}}\Big{)}$ (5)
for all $i\geq 0$ by Proposition 3.2.2. Note that this proposition can be applied to any $t_{1}\ldots t_{r}\in S^{*}$ by setting $w_{j}=t_{j}$ for $j=1,\ldots r$.
Using the description of the submonoid of reduced elements in $B^{+}$ given by Matsumoto’s theorem [GeckPfeiffer, Theorem 1.2.2, Exercise 4.1], we know that the criterion for $u$ to come from a reduced expression in the Weyl group is that $u$ has no subexpression of the form $ss$ for any $s\in S$. If $u=u_{1}ssu_{2}$, we apply Lemma 4.1.1 and Proposition 4.1.4 to obtain $u_{1}su_{2}$ such that ${\overline{X}^{S^{*}}(u)}$ and ${\overline{X}^{S^{*}}(u_{1}su_{2})}$ have the same cohomology for the structure sheaf. Inductively, this yields an element $v\in S^{*}$ with ${{\ell^{S^{*}}}(v)}\leq{{\ell^{S^{*}}}(u)}$ such that $v$ has no subexpression of the form $ss$ for any $s\in S$ and for all $i\geq 0$,
@@
By [GeckKimPfeifferMinlength, Theorem 2.6], there exists $v^{\prime}\in W$ which is of minimal length in its $F$-conjugacy class such that $\alpha(v)\to_{F}v^{\prime}$ and $\ell(v^{\prime})\leq\ell(\alpha(v))$. Suppose $\alpha(v)$ is a minimal length element in its $F$-conjugacy class. By Matsumoto’s theorem [GeckPfeiffer, Theorem 1.2.2], there exists a lift $y\in B^{+}$ of $\alpha(v)$ such that any reduced expression $x\in S^{*}$ of $\alpha(v)$ satisfies $\beta(x)=y$. The statement of the theorem thus follows from Proposition 3.2.2.
@@
In general, $\alpha(v)\to_{F}v^{\prime}$ gives $v_{1}^{\prime},\ldots,v_{m}^{\prime}\in W,s_{1},\ldots s_{m-1}\in S$ such that $v_{1}^{\prime}=\alpha(v),v_{m}^{\prime}=v^{\prime}$ and $v_{j}^{\prime}=s_{j-1}v_{j-1}^{\prime}F(s_{j-1})$ for all $j=2,\ldots,m$. Let $a\in\{2,\ldots,m\}$ be the smallest integer such that $\ell(v_{a}^{\prime})<\ell(v_{a-1}^{\prime})$. For all $j=1,\ldots,a$, let $v_{j}\in S^{*}$ be the canonical lift of the corresponding reduced expression of $v_{j}^{\prime}$. Up to taking different lifts of braid relations, for each $j=2,\ldots,a-1$, we have $v_{j}=s_{j-1}z_{j-1}$ (resp. $v_{j}=z_{j-1}F(s_{j-1})$) and $v_{j-1}=z_{j-1}F(s_{j-1})$ (resp. $v_{j-1}=s_{j-1}z_{j-1}$) for some $z_{j-1}\in S^{*}$ (cf. [GeckKimPfeifferMinlength, Remark 2.3]). Thus it follows from Corollary 4.1.5 that ${\overline{X}^{S^{*}}(v_{j})}$ all have the same cohomology for the structure sheaf as ${\overline{X}^{S^{*}}(v)}$ for $j=2,\ldots,a-1$.
@@
On the other hand, we have $v_{a-1}=s_{a-1}v_{a}F(s_{a-1})$ in $S^{*}$. We know that ${\overline{X}^{S^{*}}(v)}$ has the same cohomology as ${\overline{X}^{S^{*}}(s_{a-1}v_{a})}$ and ${\overline{X}^{S^{*}}(v_{a}F(s_{a-1}))}$ by Proposition 4.1.4. In this case, we reset $v$ to be $s_{a-1}v_{a}$. This process will terminate and yields an element $x_{0}\in W$ which is of minimal length in its $F$-conjugacy class, cf. [GeckKimPfeifferMinlength, Theorem 2.6].
@@
It remains to consider the case when $C$ is an elliptic conjugacy class in $W$. By [HeNieminlength, Corollary 4.4], we have $x\to_{F}y$ for any two minimal length elements $x,y\in C$. The corresponding cohomology of ${\overline{X}^{S^{*}}(x)}$ and ${\overline{X}^{S^{*}}(y)}$ for the structure sheaf are isomorphic by Corollary 4.1.5. ∎
Corollary 5.1.2.
Assume $F(S)=S$. Let $v_{1},\ldots,v_{r}\in W$. There exists an $F$-conjugacy class $C$ in $W$ and a minimal length element $x_{0}\in C$ such that for any reduced expression $x\in S^{*}$ of $x_{0}$, there are $G^{F}$-equivariant isomorphisms
$H^{i}\Big{(}X(\underline{v_{1}},\ldots,\underline{v_{r}}),\mathcal{O}_{X(\underline{v_{1}},\ldots,\underline{v_{r}})}\Big{)}\overset{\sim}{\longrightarrow}H^{i}\Big{(}{\overline{X}^{S^{*}}(x)},\mathcal{O}_{{\overline{X}^{S^{*}}(x)}}\Big{)}$
and
$H^{i}\Big{(}X(\underline{v_{1}},\ldots,\underline{v_{r}}),\omega_{X(\underline{v_{1}},\ldots,\underline{v_{r}})}\Big{)}\overset{\sim}{\longrightarrow}H^{i}\Big{(}{\overline{X}^{S^{*}}(x)},\omega_{{\overline{X}^{S^{*}}(x)}}\Big{)}$
for all $i\geq 0$.
Proof.
Fix reduced expressions for $v_{1},\ldots,v_{r}\in W$, and take lifts $v_{1}^{\prime},\ldots,v_{r}^{\prime}\in S^{*}$ corresponding to these reduced expressions. By Theorem 5.1.1, it suffices to show that the cohomology of $X(\underline{v_{1}},\ldots,\underline{v_{r}})$ for the structure sheaf is given by the cohomology of ${\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})}$.
Via projection, there is a proper birational morphism $f:{\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})}\to X(\underline{v_{1}},\ldots,\underline{v_{r}})$ of projective and strongly $F_{p}$-regular schemes (Proposition 2.4.1). Thus there is a $G^{F}$-equivariant isomorphism
$f_{*}\mathcal{O}_{{\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})}}\overset{\sim}{\longrightarrow}\mathcal{O}_{X(\underline{v_{1}},\ldots,\underline{v_{r}})},$
by the proof of Zariski’s main theorem in [Hartshorne, Corollary III.11.4].
The definition of local rings being strongly $F_{p}$-regular implies that they are $F_{p}$-rational. Smith has shown that $F_{p}$-rational implies pseudo-rational [Ksmith97, Theorem 3.1]. Moreover, recall that compactified Deligne–Lusztig varieties are normal and Cohen–Macaulay (Proposition 2.4.1). By [Kovacs, Theorem 8.13], the higher direct image $R^{j}f_{*}\mathcal{O}_{{\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})}}$ for $j>0$ vanish.
Functoriality of the associated Leray spectral sequence yields $G^{F}$-equivariant isomorphisms on cohomology groups
$H^{i}\Big{(}X(\underline{v_{1}},\ldots,\underline{v_{r}}),\mathcal{O}_{X(\underline{v_{1}},\ldots,\underline{v_{r}})}\Big{)}\overset{\sim}{\longrightarrow}H^{i}\Big{(}{\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})},\mathcal{O}_{{\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})}}\Big{)},$
for all $i\geq 0$. The $G^{F}$-equivariant isomorphisms for the cohomology of the canonical sheaf follow from Serre duality for Deligne–Mumford stacks [Nironi, Corollary 2.10].
@@
We use the same notations as in Theorem 5.1.1 and set $I={{\text{supp}}}_{F}(x)$. There are $G^{F}$-equivariant morphisms given by [DMR, Proposition 2.3.8]:

### v3

@@ 4.1. $\mathbb{P}^{1}$-fibrations over Deligne–Lusztig varieties
Let $w_{1},\ldots,w_{r},w_{1}^{\prime},\ldots,w_{m}^{\prime}\in W$ and $s\in S$. The projection map
@@
given by $(B_{0},\ldots,B_{r},B_{r+1},B_{r+2},\ldots,B_{r+m+1},FB_{0})\mapsto(B_{0},\ldots,B_{r},B_{r+2},\ldots,B_{r+m+1},FB_{0})$ is a $G^{F}$-equivariant Zariski $\mathbb{P}^{1}$-bundle.
The $G^{F}$-equivariant morphism of $k$-schemes
@@
For any $t\in S$, the projection map $O(\underline{t},\underline{t})\to O(\underline{t})$ given by $(B_{0},B_{1},B_{2})\mapsto(B_{0},B_{2})$ is a $\mathbb{P}^{1}$-bundle. Recall that it is a property of the Bruhat decomposition that if $(B_{0},B_{1})\in O(t)$ and $(B_{1},B_{2})\in O(t)$ then $(B_{0},B_{2})\in O(\underline{t})$, cf. [7, §1.2].
@@
Consider $t=F(s)\in S$. One can verify that $\pi$ is $G^{F}$-equivariant by its definition. Let $(B_{0},B_{1})\in O(\underline{s})$, so $(FB_{0},FB_{1})\in O(\underline{F(s)})$. Recall that via the projection to the first $(r+2)$-entries, $X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$ can be identified with a closed subscheme of $(G/B)^{r+2}$. Under this identification, we have an embedding
@@
Let $w_{1},\ldots,w_{r}\in W$ and $s\in S$ such that $\overline{Bw_{n}B}$ is smooth for $n=1,\ldots,r$. The morphisms
@@
Given $(B_{0},B_{1})\in O(\underline{s})$, we know that $(FB_{0},FB_{1})\in O(F(\underline{s}))$. By construction we have $(B_{r+1},FB_{0})\in O(F(\underline{s}))$, so $(B_{r+1},FB_{1})\in O(F(\underline{s}))$ and $(FB_{r+1},F^{2}B_{1})\in O(F(\underline{s}))$ [7, §1.2]. Thus $\tau_{1}$ and $\tau_{2}$ are well-defined. They are $G^{F}$-equivariant by construction.
@@
[…]
@@
and that $\iota$ is surjective. By the Bruhat decomposition, for every $i=1,\ldots,r$, there exists $v_{i}\in W$ such that $(B_{i},B_{i}^{\prime})\in O(v_{i})$. This implies $(FB_{i},FB_{i}^{\prime})\in O(F(v_{i}))$. We know $FB_{i}=FB_{i}^{\prime}$ by assumption, so $F(v_{i})=e$ is the neutral element in $W$ and $v_{i}=e$ because $F:W\to W$ is an automorphism, cf. [7, §1.2]. It follows that $\iota$ is also injective and an isomorphism.
Hence $d_{1}$ is a Zariski $\mathbb{P}^{1}$-bundle. Since $F:X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})\to X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})$ is faithfully flat by our assumption on smoothness [13, Corollary 14.129], we deduce that $\tau_{2}$ is faithfully flat by faithfully flat descent [13, Corollary 14.12, Proposition 14.3(2)].
@@
Let $w_{1},\ldots,w_{r}\in W$ and $s\in S$ such that $\overline{Bw_{n}B}$ is smooth for $n=1,\ldots,r$. There are $G^{F}$-equivariant isomorphisms for all $i\geq 0$,
@@
The proof is simpler for the Zariski $\mathbb{P}^{1}$-bundles $\phi$ and $\pi$, so we only record the proof for the fppf $\mathbb{P}^{1}$-bundles $\tau_{1}$ and $\tau_{2}$. We use the notations of Lemma 4.1.3. Since $d_{1}$ is a Zariski $\mathbb{P}^{1}$-bundle, it is proper and faithfully flat. Hence $\tau_{1}$ and $\tau_{2}$ are proper and faithfully flat by faithfully flat descent [13, Proposition 14.53].
@@
Let $x\in X_{2}$ and $y\in X_{2}^{\prime}$ such that $F(y)=x$ as in diagram (2). By Lemma 4.1.3, the base change of the fibre $x\times_{X_{2}}X_{1}$ to $y$ is isomorphic to $\mathbb{P}^{1}_{y}$. It follows from flat base change that $H^{j}(x\times_{X_{2}}X_{1},\mathcal{O}_{x\times_{X_{2}}X_{1}})=0$ for all $j\geq 0$ [14, Corollary 22.91]. We conclude by Grauert’s theorem [14, Proposition 23.142] (cf. [13, Proposition 3.35]) that $R^{j}f_{*}\mathcal{O}_{X_{1}}=0$ for all $j>0$. Hence the Leray spectral sequence degenerates.
We have an fppf covering $\mathfrak{U}=\{\gamma_{\alpha}:U_{\alpha}\to X_{2}\}_{\alpha}$ of $X_{2}$ in Lemma 4.1.3, such that the base change $f_{\alpha}:X_{1}\times_{X_{2}}U_{\alpha}\to U_{\alpha}$ of $f$ along $\gamma_{\alpha}$ is a trivial $\mathbb{P}^{1}$-bundle for all $\alpha$. Since $U_{\alpha}$ is affine and $\gamma_{\alpha}$ is an affine morphism for all $\alpha$ [13, Proposition 12.3(3)], any simplex $V$ in the Čech nerve of $\mathfrak{U}$ is affine and for any coherent sheaf $\mathcal{F}$ on $X_{2}$, we have $H^{i}(X_{2},\mathcal{F})\overset{\sim}{\longrightarrow}\check{H}^{i}(\mathfrak{U},\mathcal{F})$ for all $i\geq 0$ [12, Theorem II.5.4.1, Corollary]. Note that $f_{*}\mathcal{O}_{X_{1}}$ is a coherent $\mathcal{O}_{X_{2}}$-module because $f$ is proper and cohomology of coherent sheaves is independent of topology [2, 03DW].
For each simplex $V$ in the Čech nerve of $\mathfrak{U}$, we have $V\times_{X_{2}}X_{1}\overset{\sim}{\longrightarrow}V\times_{{\rm Spec}\ k}\mathbb{P}^{1}$, so $f_{*}\mathcal{O}_{X_{1}}(V)=\mathcal{O}_{V}(V)$ by flat base change [14, Proposition 22.90]. This identification is compatible with restriction of sections, so the Čech complexes on $\mathfrak{U}$ for $f_{*}\mathcal{O}_{X_{1}}$ and $\mathcal{O}_{X_{2}}$ are identical and $H^{i}(X_{2},\mathcal{O}_{X_{2}})\overset{\sim}{\longrightarrow}H^{i}(X_{2},f_{*}\mathcal{O}_{X_{1}})$ for all $i\geq 0$.
@@
$H^{i}\left(X_{2},\mathcal{O}_{X_{2}}\right)\overset{\sim}{\longrightarrow}H^{i}\left(X_{1},\mathcal{O}_{X_{1}}\right),$
@@
The $G^{F}$-equivariant isomorphisms for the cohomology of $\omega_{X_{1}}$ and $\omega_{X_{2}}$ follow from Serre duality for Deligne–Mumford stacks [24, Corollary 2.10].
@@
Let $s\in S$. Let $w=sv$ and $w^{\prime}=vF(s)$ be elements of $S^{*}$. There are $G^{F}$-equivariant isomorphisms for all $i\geq 0$,
@@ 4.2. $F$-conjugacy classes in $W$
Minimal length elements were studied by Geck, Kim, and Pfeiffer [10], as well as He and Nie [19]. In our context, the results of these two papers imply that that for any $w\in W$, there exists a minimal length element $w^{\prime}$ in the $F$-conjugacy class of $w$ such that $w\to_{F}w^{\prime}$ (cf. Definition 4.2.1). In combination with Proposition 4.1.4, we obtain a framework for an inductive approach for studying cohomology groups of Deligne–Lusztig varieties.
@@
Furthermore, for any pairs of minimal length elements $w$ and $w^{\prime}$ in an elliptic $F$-conjugacy class $C$ in $W$, we have $w\to_{F}w^{\prime}$ and $w^{\prime}\to_{F}w$, cf. [19, Corollary 4.4]. The definition of an elliptic $F$-conjugacy class, which we do not recall here, can be found in §4.1 of loc. cit.
@@ 5.1. Main theorem and applications
For each $w\in S^{*}$, there exists an $F$-conjugacy class $C$ in $W$ and a minimal length element $x_{0}\in C$ such that for any reduced expression $x\in S^{*}$ of $x_{0}$, there are $G^{F}$-equivariant isomorphisms
@@
We prove the theorem for the structure sheaf. The results for the canonical sheaf follows from Serre duality for Deligne–Mumford stacks [24, Corollary 2.10].
Let $w\in S^{*}$. We showed in Proposition 3.2.1 that the cohomology for the structure sheaf of $\overline{X}^{S^{*}}(w)$} and $\overline{X}^{S^{*}}(w^{\prime})$} are isomorphic for any other lift $w^{\prime}$ of $\beta(w)$ under the projection $\beta:S^{*}\twoheadrightarrow B^{+}$.
Using the description of the submonoid of reduced elements in $B^{+}$ given by Matsumoto’s theorem [11, Theorem 1.2.2, Exercise 4.1], we know that the criterion for $u\in B^{+}$ to come from a reduced expression in the Weyl group is that $u$ has no subexpression of the form $ss$ for any $s\in S$.
If $w=w_{1}ssw_{2}$, we apply Lemma 4.1.1 and Proposition 4.1.4 to obtain $w_{1}sw_{2}$ such that $\overline{X}^{S^{*}}(w)$} and $\overline{X}^{S^{*}}(w_{1}sw_{2})$} have the same cohomology for the structure sheaf. Inductively, this yields an element $v\in S^{*}$ with ${{\ell^{S^{*}}}(v)}\leq{{\ell^{S^{*}}}(w)}$ such that $v$ has no subexpression of the form $ss$ for any $s\in S$ and for all $i\geq 0$,
@@
By [10, Theorem 2.6], there exists $v^{\prime}\in W$ which is of minimal length in its $F$-conjugacy class such that $\alpha(v)\to_{F}v^{\prime}$ and $\ell(v^{\prime})\leq\ell(\alpha(v))$. Suppose $\alpha(v)$ is a minimal length element in its $F$-conjugacy class. By Matsumoto’s theorem [11, Theorem 1.2.2], there exists a lift $y\in B^{+}$ of $\alpha(v)$ such that any reduced expression $x\in S^{*}$ of $\alpha(v)$ satisfies $\beta(x)=y$. The statement of the theorem thus follows from Proposition 3.2.1.
@@
In general, $\alpha(v)\to_{F}v^{\prime}$ gives $v_{1}^{\prime},\ldots,v_{m}^{\prime}\in W,s_{1},\ldots s_{m-1}\in S$ such that $v_{1}^{\prime}=\alpha(v),v_{m}^{\prime}=v^{\prime}$ and $v_{j}^{\prime}=s_{j-1}v_{j-1}^{\prime}F(s_{j-1})$ for all $j=2,\ldots,m$. Let $a\in\{2,\ldots,m\}$ be the smallest integer such that $\ell(v_{a}^{\prime})<\ell(v_{a-1}^{\prime})$. For all $j=1,\ldots,a$, let $v_{j}\in S^{*}$ be the canonical lift of the corresponding reduced expression of $v_{j}^{\prime}$. Up to taking different lifts of braid relations, for each $j=2,\ldots,a-1$, we have $v_{j}=s_{j-1}z_{j-1}$ (resp. $v_{j}=z_{j-1}F(s_{j-1})$) and $v_{j-1}=z_{j-1}F(s_{j-1})$ (resp. $v_{j-1}=s_{j-1}z_{j-1}$) for some $z_{j-1}\in S^{*}$ (cf. [10, Remark 2.3]). Thus it follows from Corollary 4.1.5 that $\overline{X}^{S^{*}}(v_{j})$} all have the same cohomology for the structure sheaf as $\overline{X}^{S^{*}}(v)$} for $j=2,\ldots,a-1$.
@@
On the other hand, we have $v_{a-1}=s_{a-1}v_{a}F(s_{a-1})$ in $S^{*}$. We know that $\overline{X}^{S^{*}}(v)$} has the same cohomology as $\overline{X}^{S^{*}}(s_{a-1}v_{a})$} and $\overline{X}^{S^{*}}(v_{a}F(s_{a-1}))$} by Proposition 4.1.4. In this case, we reset $v$ to be $s_{a-1}v_{a}$. This process will terminate and yields an element $x_{0}\in W$ which is of minimal length in its $F$-conjugacy class, cf. [10, Theorem 2.6].
@@
It remains to consider the case when $C$ is an elliptic conjugacy class in $W$. By [19, Corollary 4.4], we have $x\to_{F}y$ for any two minimal length elements $x,y\in C$. The corresponding cohomology of $\overline{X}^{S^{*}}(x)$} and $\overline{X}^{S^{*}}(y)$} for the structure sheaf are isomorphic by Corollary 4.1.5. ∎
@@
We use the same notations as in Theorem˜5.1.1 and set $I={{\text{supp}}}_{F}(x)$. There are $G^{F}$-equivariant morphisms given by [8, Proposition 2.3.8]: